OpenStudy (anonymous):

describe the motion of a point on a wheel that has a centre 4m off the ground, has radius of 15 cm, makes a full rotation every 10 seconds and starts at its highest point. use y=asink(t-b)+c model or cosine function.

OpenStudy (anonymous):

c=4 and a=.15 ... is that part clear?

OpenStudy (anonymous):

a=15 not .15..

OpenStudy (anonymous):

false. :) continuing... \[f=\frac{ 1 }{10}\] \[\omega = 2 \pi f\] \[w= \frac{ \pi }{ 5 }\]

OpenStudy (anonymous):

agree so far?

OpenStudy (anonymous):

yes i got that part

OpenStudy (anonymous):

so the phase...

OpenStudy (anonymous):

we want it to be at the highest point when t=0... sin =1 when theta = pi/2 pi/5 *(t-b) = pi/2 pi/5(-b) = pi/2 b= 5/2

OpenStudy (anonymous):

\[y = .15 \sin (\frac{ \pi }{ 5} (t+\frac{ 5 }{2})) +4\]

OpenStudy (anonymous):

would it be appropriate to keep(t-b) it at pi/5 *(t) + pi/2 ?

OpenStudy (anonymous):

you can... that's not the form they ask you to use though...

OpenStudy (anonymous):

whoops, should say b=-5/2 up there... typo

OpenStudy (anonymous):

oh okay. & also for cosine model it would just be \[y=.15\cos(\frac{ \pi }{ 5 }t) +4\] right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why would it be -5/2 and not 5/2 ? :S

OpenStudy (anonymous):

why is b= -5/2? is that what you're asking?

OpenStudy (anonymous):

OH WAIT.. nvm.. you were referring to your reply before the equation. nvm. lol

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

sure:)